3x + 6x-9 -6 = 0 for x
over x-1 over x-1
3x over x-1 plus 6x-9 over x-1 minus 6 equals 0 solve for x?

Answer by checkley77(12844)

3x/(x-1)+(6x-9)/(x-1)-6=0
(3x+6x-9)/(x-1)=6
3x+6x-9=6(x-1)
9x-9=6x-6
9x-6x=-6+9
3x=3
x=3/3
x=1 answer.
proof:
3*1/(1-1)+(6*1-9)/(1-1)-6=0
3/0+3/0-6=0 this is an invalid process by dividing by 0 thus there is no valid solution.